# Maximizing Battery Storage Profits via High-Frequency Intraday Trading

### v2

@@ Abstract
Maximizing revenue for grid-scale battery energy storage systems in continuous intraday electricity markets requires strategies that are able to seize trading opportunities as soon as new information arrives. This paper introduces and evaluates an automated high-frequency trading strategy for battery energy storage systems trading on the intraday market for power while explicitly considering the dynamics of the limit order book, market rules, and technical parameters. The standard rolling intrinsic strategy is adapted for continuous intraday electricity markets and solved using a dynamic programming approximation that is two to three orders of magnitude faster than an exact mixed-integer linear programming solution. A detailed backtest over a full year of German order book data demonstrates that the proposed dynamic programming formulation does not reduce trading profits and enables the policy to react to every relevant order book update, enabling realistic rapid backtesting. Our results show the significant revenue potential of high-frequency trading: our policy earns $58$% more than when re-optimizing only once every hour and $14$% more than when re-optimizing once per minute, highlighting that profits critically depend on trading speed. Furthermore, we leverage the speed of our algorithm to train a parametric extension of the rolling intrinsic, increasing yearly revenue by $8.4$% out of sample.
@@ 1 Introduction
    We parametrize the standard formulation of the RI, and show that our simulation speed can be used to easily find optimal parameters, generating additional profits of $8.5$, by slightly nudging the RI’s trading behavior at no additional risk.
@@ 2 The European Power Market
The European electricity markets operate primarily within two major trading regions, each with their own dedicated exchange: the NordPool exchange, covering the Scandinavian region, and the EPEX exchange, which serves most of the rest of Europe.
@@ 2.1 The Intraday Market for Power
Figure 1: An exemplary state of the LOB at $T=1$ presented in the left panel. The right panel depicts the clearing of a new hypothetical buy order: A new buy order with a price of 50 €, which is higher than the lowest ask price, is added to the book at $T=2$. The quantity of this buy order (30 MW) is then cleared against the cheapest possible offers until either the whole order is fulfilled (as is the case in the figure) or until there are no offers with lower prices left. In this example, $22$ MW out of $30$ MW are cleared against the sell order with price 37 € and the remaining 30 MW $-\ 22$ MW $=8$ MW are cleared against the sell order with price 45 €. The remaining quantity of 12 MW $-\ 8$ MW $=4$ MW of the latter order stays in the order book. Note that the clearing is instantaneous, i.e., columns 2–4 in the right panel are purely illustrative and do not correspond to market states that can be observed by traders. Figure and caption adapted from Graf et al. (2024).
@@ 3.1 The Intrinsic Problem
The intrinsic strategy is a simple policy to maximize the rewards of a storage system, which is widely used in the literature on gas storage and in the industry to optimize storage assets (e.g. Gray and Khandelwal, 2004b, a; Lai et al., 2010, 2011; Löhndorf and Wozabal, 2021). In the extant literature, the intrinsic problem is set in a perfectly liquid futures market, trading $T\in\mathds{N}$ futures of a commodity that can be stored and which trade for prices $P_{t}$. In order not to overcomplicate notation, we assume here that future contracts are for delivery in consecutive time periods $[0,1),[1,2),\dots,[T-1,T)$ and that there are no overlapping contracts. To fix ideas, one can imagine a battery storage system trading futures for delivery in the $24$ hours of the next day.
@@
where the $\alpha_{t}$ are binary variables which are equal to $1$ in case the decision for contract $t$ is to buy and 00 if the decision is to sell.
@@ 3.3 A Dynamic Programming Formulation of the Rolling Intrinsic Policy
We propose solving the problem by an application of the dynamic programming algorithm, discretizing the actions $f_{t}$ into steps $v\cdot u$ ($v\in\mathds{N}$) apart, where $u$ is the minimal tradable quantity. Furthermore, we introduce a function $\pi_{t}$ that encodes the current order book information and returns the cost or revenue of buying or selling on the intraday market, taking into account trading costs. Furthermore, we define a state transition function
@@
Therefore, unlike classic applications of the dynamic programming algorithm, we do not define $V_{T}$ to have a finite domain, but instead on the whole interval of possible storage states $[0,\bar{s}]$. In order to obtain the value function, we discretize the possible storage states $[0,\bar{s}]$ to a finite grid $G\subset[0,\bar{s}]$ and $\lvert G\rvert=m$. We then approximate the value function $V_{T}$ by evaluating it on the grid and linearly interpolating for points $s\notin G$.
@@
Input: $f_{t}^{0}$, $\mathcal{O}_{t}$ for all tradable products $t\in\mathcal{T}$, other parameters $(v,s_{0},\nu,G,\dots)$
@@
Suppose that the initial state of charge (SoC) is $s_{0}$, we set $v=1$ and look at the simplest case of $f_{t}^{0}=0$ for all $t\in\mathcal{T}$. The possible storage states at the beginning of stage $t_{0}+1$ are
@@
$\mathcal{K}_{t}^{+}$ =$=\left\{k\in\mathds{N}_{0}:k\leq\frac{(\bar{s}-s_{0})+u\max(\mathcal{K}_{t-1}^{-})(\eta^{-})^{-1}}{u\eta^{+}}\right\}$
$\mathcal{K}_{t}^{-}$ =$=\left\{k\in\mathds{N}_{0}:k\leq u^{-1}\eta^{-}\left(s_{0}+\eta^{+}u\max(\mathcal{K}_{t-1}^{+})\right)\right\}.$
@@ 4.1 Setting and Mode of Comparison
We use linear degradation costs of $\nu_{\text{deg}}=4$/MWh, i.e., a round-trip degradation cost of $8$/MWh for every MWh of electricity sold by the battery. We motivate this value by Goldman Sachs (2023) who project the cost of battery replacement in 2030 to be around $60$k €/ MWh, a 10-year battery warranty, and approximately $2$ charging cycles per day. Dividing the replacement costs by cycles gives the cost of one full cycle, which allows us to infer the induced per MWh degradation cost.
Furthermore, we choose $\nu_{\text{trade}}=0.09$/MWh, reflecting the current EPEX volume-based trading fees per each matched order. In total, the variable cost $\nu=\nu_{\text{trade}}+\nu_{\text{deg}}$ thus takes the value of $\nu=4.09$/MWh.
All results are generated using historical 2021 order book data from the EPEX Spot continuous intraday market for Germany, focusing on hourly products. The exchange allows for price increments of $0.01$ and volume increments at $0.1$ MWh, defining the minimum action size on the market as $u=0.1$ MWh. The discretization of the dynamic programming action is defined at the most accurate minimum value of $v=1$ throughout.
@@
If the combined delay exceeds $1$ ms, this has two effects on trading: Firstly, there is the risk that an order placed by our policy may not be matched since its intended counterpart either expired or was cleared against other market participants’ orders in the meantime. We deal with this by submitting all-or-none orders that expire if they are not immediately cleared, which might lead to non-physical positions, i.e., market positions that cannot be matched by a feasible schedule of injections and withdrawals. In such a situation, further trades are required to correct the position so that the final battery schedule becomes physically feasible. However, in the way the intrinsic problem is set up, such positions would be corrected in the next call of the intrinsic problem, forcing the RI to return to a physical position as quickly as possible.
@@
Finally, we remark that when backtesting, we correctly account for changes in the order book caused by our actions. This in particular means that orders cleared against the orders of the policy are available shorter than in the actual history, and consequently that subsequent clearing of orders by other markets participants changes.
@@
All simulations were implemented in C++ and run on a single AMD EPYC 9654 CPU, with the MILP solved using Gurobi v11.0.3 for C++ (Gurobi Optimization, LLC, 2024). Due to the simple nature of each single intrinsic optimization, solving the MILP on multiple threads does not lead to faster simulation times. We publish or code in the Python package BitePy, which allows users to easily run our DP simulations over a Python interface. See A for details.
@@ 4.2 Comparison between MILP and DP Rolling Intrinsic
We choose the most clinical setting for our comparison, by artificially fixing the solve time and technical delay to 00 ms for both approaches. This results in a policy that ignores technical delays and solves the intrinsic optimization for each relevant order book update in such a way that the resulting all-or-none limit orders are matched instantaneously with the exchange after each solve.
@@ 4.3 Yearly High-Frequency Rolling Intrinsic Trading
Leveraging the speed of our proposed DP method, we present full-year RI trading results under a standard battery and simulation setup for the year 2021 in Figure 3. This figure compares rewards, of RI settings outlined in Section 4.1 with a value function discretization setting of $m=11$ and a technical delay of $200$ ms plus measured actual solve times. We evaluate performance starting from the baseline with a solve-frequency of $60$ minutes, i.e. solving the intrinsic optimization once every hour, resulting in a reward of €$221$k (8.7k orders submitted at the exchange) and a simulation runtime of $5.6$ minutes, down to solving it with every relevant order book update (apart from the waiting durations introduced previously) with a reward of €$349$k (30k orders submitted at the exchange) and a simulation runtime $86$ minutes. We define any order message submitted to the LOB as relevant, if it is placed at the head of its corresponding bid/ask stack. These results clearly demonstrate that solving the intrinsic at higher frequencies leads to progressively increasing profits as the trading frequency rises. In this case, a $58\%$ increase between our slowest and quickest solve frequency. We observe a clear increase in the slope of the cumulative reward starting in October 2021. This originates most likely from higher and more volatile gas prices in Europe during this time, which subsequently translates to more extreme intraday electricity prices, which increases the revenue potential of storage.
@@ 4.4 Parameter Evaluation
Additionally, batteries with a smaller maximum storage capacity produce larger profits per MW storage capacity, as the increased trading flexibility of larger batteries profits from the same most profitable trading opportunities as with a smaller battery, and only adds less profitable trades on top. For similar reasons, a battery with a duration of $2$h yields lower profits, but more than half of the profits observed for our standard setup with a duration of $1$h.
@@ 4.5 Parametrization of the Rolling Intrinsic
The increased speed of our DP solution method allows us to quickly search for optimal parameters $\phi$. For the sake of simplicity and just to show a proof of concept, we fix the parameters $\phi$ at constant values for each trading month, and use the previous month $m-1$ to train $\phi$ for a given month $m$. When evaluating this strategy over the entire year 2021, this sliding window parameter training results in $12$ separate training windows. We train the parameters on the highest intrinsic-solve frequency (see section 4.3) using DP and simulation settings outlined in Section 4.1, which represents the most realistic training scenario. We treat our monthly simulation rewards as a black box function $f(\phi)$ and, given its simple dependence on a singular parameter, use Brent’s method to find its maximum.
The resulting optimal parameters for the year 2021 have an average penalty of $\bar{\phi}=2.1\pm 1.6$. Comparing year-long out-of-sample results of the newly trained policy yields an increase of $8.4$% in trading reward over the entire year 2021, accumulating €$376$k, compared to the €$347$k reward of the standard policy with the penalty set to 00, i.e., $\phi\equiv 0$.
@@ 5 Conclusion
An important future extension of our work lies in more sophisticated parametrizations of the rolling intrinsic. Our naive method could easily be adapted to multiple parameters, affecting various aspects of the RI’s trading behavior, and utilizing the speed of our optimization method. Furthermore, future work could naturally extend the trading to a multi-market scenario, where additional decision variables expand the scope of optimization to multiple subsequent electricity markets, e.g. the reserve or day-ahead markets. Finally, extending our intrinsic optimization with dynamic price forecasts would greatly improve the RI’s profitability, as the optimization would be able to anticipate future price changes. Fine-tuning forecasts to BESS applications would then be a logical utilization of the speed of our method.

### v3

@@ Abstract
Maximizing revenue for grid-scale battery energy storage systems in continuous intraday electricity markets requires strategies that are able to seize trading opportunities as soon as new information arrives. This paper introduces and evaluates an automated high-frequency trading strategy for battery energy storage systems trading on the intraday market for power while explicitly considering the dynamics of the limit order book, market rules, and technical parameters. The standard rolling intrinsic strategy is adapted for continuous intraday electricity markets and solved using a dynamic programming approximation that is two to three orders of magnitude faster than an exact mixed-integer linear programming solution. A detailed backtest over a full year of German order book data demonstrates that the proposed dynamic programming formulation does not reduce trading profits and enables the policy to react to every relevant order book update, enabling realistic rapid backtesting. Our results show the significant revenue potential of high-frequency trading: our policy earns 5858% more than when re-optimizing only once every hour and 1414% more than when re-optimizing once per minute, highlighting that profits critically depend on trading speed. Furthermore, we leverage the speed of our algorithm to train a parametric extension of the rolling intrinsic, increasing yearly revenue by 8.48.4% out of sample.
@@ 1 Introduction
    We parametrize the standard formulation of the RI, and show that our simulation speed can be used to easily find optimal parameters, generating additional profits of $8.5$ %, by slightly nudging the RI’s trading behavior at no additional risk.
@@ 2 The European Power Market
The European short-term electricity markets are served by two primary competing power exchanges, Nord Pool and EPEX SPOT. While Nord Pool originated in the Nordic region, both exchanges now operate across wide areas of Europe, including the Nordic and Central Western European (CWE) regions.
@@ 2.1 The Intraday Market for Power
[…]
Figure 1: An exemplary state of the LOB at $T=1$ presented in the left panel. The right panel depicts the clearing of a new hypothetical buy order: A new buy order with a price of 50 €/MWh, which is higher than the lowest ask price, is added to the book at $T=2$. The quantity of this buy order (30 MWh) is then cleared against the cheapest possible offers until either the whole order is fulfilled (as is the case in the figure) or until there are no offers with lower prices left. In this example, $22$ MWh out of $30$ MWh are cleared against the sell order with price 37 €/MWh  and the remaining 30 MWh 22-\ 22 MWh =8=8 MWh are cleared against the sell order with price 45 €/MWh. The remaining quantity of 12 MWh 8-\ 8 MWh =4=4 MW of the latter order stays in the order book. Note that the clearing is instantaneous, i.e., columns 2–4 in the right panel are purely illustrative and do not correspond to market states that can be observed by traders. Figure and caption adapted from Graf et al. (2024).
@@ 3.1 The Intrinsic Problem
The intrinsic strategy is a simple policy to maximize the rewards of a storage system, which is widely used in the literature on gas storage and in the industry to optimize storage assets (e.g. Gray and Khandelwal, 2004b, a; Lai et al., 2010, 2011; Löhndorf and Wozabal, 2021). In the extant literature, the intrinsic problem is set in a perfectly liquid futures market, trading $T\in\mathds{N}$ futures of a commodity that can be stored and which trade for prices $P_{t}$. In order not to overcomplicate notation, we assume here that future contracts are for delivery in consecutive time periods $0,1),[1,2),\dots,[T-1,T)$ and that there are no overlapping contracts. To fix ideas, one can imagine a battery storage system trading futures for delivery in the $24$ hours of the next day.
@@
where the $\alpha_{t}$ are binary variables which are equal to $1$ in case the decision for contract $t$ is to buy and 0 if the decision is to sell.
@@ 3.3 A Dynamic Programming Formulation of the Rolling Intrinsic Policy
We propose solving the problem by an application of the dynamic programming algorithm, discretizing the actions $f_{t}$ into steps $\kappa\cdot u$ ($\kappa\in\mathds{N}$) apart, where $u$ is the minimal tradable quantity. Furthermore, we introduce a function $\pi_{t}$ that encodes the current order book information and returns the cost or revenue of buying or selling on the intraday market, taking into account trading costs. Furthermore, we define a state transition function
@@
Therefore, unlike classic applications of the dynamic programming algorithm, we do not define $V_{T}$ to have a finite domain, but instead on the whole interval of possible storage states $0,\bar{s}]$. In order to obtain the value function, we discretize the possible storage states $0,\bar{s}]$ to a finite grid $G\subset[0,\bar{s}]$ and $\lvert G\rvert=m$. We then approximate the value function $V_{T}$ by evaluating it on the grid and linearly interpolating for points $s\notin G$.
@@
Input: $f_{t}^{0}$, $\mathcal{O}_{t}$ for all tradable products $t\in\mathcal{T}$, other parameters $(\kappa,s_{0},\nu,G,\dots)$
@@
Suppose that the initial state of charge (SoC) is $s_{0}$, we set $\kappa=1$ and look at the simplest case of $f_{t}^{0}=0$ for all $t\in\mathcal{T}$. The possible storage states at the beginning of stage $t_{0}+1$ are
@@
$\mathcal{K}_{t}^{+}$ $=\left\{k\in\mathds{N}_{0}:k\leq\frac{(\bar{s}-s_{0})+u\max(\mathcal{K}_{t-1}^{-})(\eta^{-})^{-1}}{u\eta^{+}}\right\}$
$\mathcal{K}_{t}^{-}$ $=\left\{k\in\mathds{N}_{0}:k\leq u^{-1}\eta^{-}\left(s_{0}+\eta^{+}u\max(\mathcal{K}_{t-1}^{+})\right)\right\}.$
@@ 4.1 Setting and Mode of Comparison
We use linear degradation costs of $\nu_{\text{deg}}=4$ €/MWh, i.e., a round-trip degradation cost of $8$ €/MWh for every MWh of electricity sold by the battery. We motivate this value by Goldman Sachs (2023) who project the cost of battery replacement in 2030 to be around 6060k €/ MWh, a 10-year battery warranty, and approximately $2$ charging cycles per day. Dividing the replacement costs by cycles gives the cost of one full cycle, which allows us to infer the induced per MWh degradation cost.
Furthermore, we choose $\nu_{\text{trade}}=0.09$ €/MWh, reflecting the current EPEX volume-based trading fees per each matched order. In total, the variable cost $\nu=\nu_{\text{trade}}+\nu_{\text{deg}}$ thus takes the value of $\nu=4.09$ €/MWh. This variable cost $\nu$ is always present in the optimization, while $\nu_{\text{trade}}$ is deducted from profits after each transaction. The degradation cost $\nu_{\text{deg}}$ is subtracted only once, after each product’s final schedule is set at gate closure.
All results are generated using historical 2021 order book data from the EPEX Spot continuous intraday market for Germany, focusing on hourly products. It is not possible ex post to simulate iceberg orders correctly given the EPEX dataset, so we omit them in our simulations. The exchange allows for price increments of $0.01$ € and volume increments at $0.1$ MWh, defining the minimum action size on the market as $u=0.1$ MWh. The discretization of the dynamic programming action is defined at the most accurate minimum value of $\kappa=1$ throughout.
@@
If the combined delay exceeds $1$ ms (the tick duration of the EPEX Spot market), this has two effects on trading: Firstly, there is the risk that an order placed by our policy may not be matched since its intended counterpart either expired or was cleared against other market participants’ orders in the meantime. We deal with this by submitting all-or-none orders that expire if they are not immediately cleared, which might lead to non-physical positions, i.e., market positions that cannot be matched by a feasible schedule of injections and withdrawals. In such a situation, further trades are required to correct the position so that the final battery schedule becomes physically feasible. However, in the way the intrinsic problem is set up, such positions would be corrected in the next call of the intrinsic problem, forcing the RI to return to a physical position as quickly as possible.
@@
Finally, we remark that when backtesting, we correctly account for changes in the order book caused by our actions. This in particular means that EPEX orders cleared against the orders of our policy are available shorter than in the actual history, and consequently that subsequent clearing of orders by other market participants changes. This introduces a strong path-dependency of varying trading policies.
@@
All simulations were implemented in C++ and run on a single AMD EPYC 9654 CPU, with the MILP solved using Gurobi v11.0.3 for C++ (Gurobi Optimization, LLC, 2024). Due to the simple nature of each single intrinsic optimization, solving the MILP on multiple threads does not lead to faster simulation times. Additionally, to provide a fair comparison, we set our our MILP solver’s time-limit to 2 seconds (MIP gap 10%) and warmstart each optimization with the null-action, i.e. the solution of not changing the current battery schedule. We publish or code in the Python package BitePy, which allows users to easily run our DP simulations over a Python interface. See A for details.
@@ 4.2 Comparison between MILP and DP Rolling Intrinsic
We choose the most clinical setting for our comparison, by artificially fixing the solve time and technical delay to 0 ms for both approaches. This results in a policy that ignores technical delays and solves the intrinsic optimization for each relevant order book update in such a way that the resulting all-or-none limit orders are matched instantaneously with the exchange after each solve.
@@ 4.3 Yearly High-Frequency Rolling Intrinsic Trading
Leveraging the speed of our proposed DP method, we present full-year RI trading results under a standard battery and simulation setup for the year 2021 in Figure 3. This figure compares rewards, of RI settings outlined in Section 4.1 with a value function discretization setting of $m=11$ and a technical delay of $200$ ms plus measured actual solve times. We evaluate performance starting from the baseline with a solve-frequency of $60$ minutes, i.e. solving the intrinsic optimization once every hour, resulting in a reward of €221221k (8.7k orders submitted at the exchange) and a simulation runtime of $5.6$ minutes, down to solving it with every relevant order book update (apart from the waiting durations introduced previously) with a reward of €349349k (30k orders submitted at the exchange) and a simulation runtime $86$ minutes. We define any order message submitted to the LOB as relevant, if it is placed at the head of its corresponding bid/ask stack. These results clearly demonstrate that solving the intrinsic at higher frequencies leads to progressively increasing profits as the trading frequency rises. In this case, a $58\%$ increase between our slowest and quickest solve frequency. We observe a clear increase in the slope of the cumulative reward starting in October 2021. This originates most likely from higher and more volatile gas prices in Europe during this time, which subsequently translates to more extreme intraday electricity prices, which increases the revenue potential of storage.
@@ 4.4 Parameter Evaluation
Additionally, batteries with a smaller maximum storage capacity produce larger profits per MW storage capacity, as the increased trading flexibility of larger batteries profits from the same most profitable trading opportunities as with a smaller battery, and only adds less profitable trades on top. For similar reasons, a battery with a duration of 22h yields lower profits, but more than half of the profits observed for our standard setup with a duration of 11h.
@@ 4.5 Parametrization of the Rolling Intrinsic
The increased speed of our DP solution method allows us to quickly search for optimal parameters $\phi$. For the sake of simplicity and just to show a proof of concept, we fix the parameters $\phi$ at constant values for each trading month, and use the previous month m1m-1 to train $\phi$ for a given month $m$. When evaluating this strategy over the entire year 2021, this sliding window parameter training results in $12$ separate training windows. We train the parameters on the highest intrinsic-solve frequency (see section 4.3) using DP and simulation settings outlined in Section 4.1, which represents the most realistic training scenario. We treat our monthly simulation rewards as a black box function $f(\phi)$ and, given its simple dependence on a singular parameter, use Brent’s method to find its maximum.
The resulting optimal parameters for the year 2021 have an average penalty of $\bar{\phi}=2.1\pm 1.6$. Comparing year-long out-of-sample results of the newly trained policy yields an increase of 8.48.4% in trading reward over the entire year 2021, accumulating €376376k, compared to the €347347k reward of the standard policy with the penalty set to 0, i.e., $\phi\equiv 0$.
@@ 5 Conclusion
An important future extension of our work lies in more sophisticated parametrizations of the rolling intrinsic. Our naive method could easily be adapted to multiple additional parameters like trading delays, affecting various aspects of the RI’s trading behavior, further utilizing the speed of our optimization method. Furthermore, future work could naturally extend the trading to a multi-market scenario, where additional decision variables expand the scope of optimization to multiple subsequent electricity markets, e.g. the reserve or day-ahead markets. Finally, extending our intrinsic optimization with dynamic price forecasts would greatly improve the RI’s profitability, as the optimization would be able to anticipate future price changes. Fine-tuning forecasts to BESS applications would then be a logical utilization of the speed of our method.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used Gemini 2.5 and ChatGPT in order to improve language and readability. After using this tool/service, the author reviewed and edited the content as needed and take full responsibility for the content of the published article.